OpenStudy (anonymous):

What is the 7th term of (3x + 2y)9?

OpenStudy (perl):



OpenStudy (anonymous):

(3x + 2y)^9?

OpenStudy (perl):

19683 x^9+118098 x^8 y+314928 x^7 y^2+489888 x^6 y^3+489888 x^5 y^4+326592 x^4 y^5+145152 x^3 y^6+41472 x^2 y^7+6912 x y^8+512 y^9

OpenStudy (anonymous):

I wanted someone to explain it too

OpenStudy (jessiegonzales):

3x=9-2y

OpenStudy (jessiegonzales):

x=9/3-2/3y

OpenStudy (jessiegonzales):

x=3-2/3y